26. A ball with a horizontal speed of 1 m/s rolls off a bench 2 m high.


How long will the ball take to reach the floor?


How far from a point on the floor directly below the edge of the bench will the ball land?



27. An electron is ejected horizontally at a speed of 1.5 X10⁶ m/s from the electron gun of a computer monitor. If the viewing screen is 35 cm from the end of the gun, how far will the electrons travel in the vertical direction before hitting the screen?



1
Expert's answer
2022-03-15T16:50:15-0400

Solution.

"26. h=2m;"

"v=1 m\/s;"

"t-?;"

"L-?;"

"h=\dfrac{gt^2}{2}\implies t=\sqrt{\dfrac{2h}{g}};"

"t=\sqrt{\dfrac{2\sdot2}{9.8}}=0.64s;"

"L=vt;"

"L=1\sdot0.64=0.64m;"

"27. v=1.5\sdot 10^6 m/s;"

"L=0.35m;"

"L=vt\implies t=\dfrac{L}{v};"

"t=\dfrac{0.35}{1.5\sdot10^6}=0.23\sdot10^{-6}s;"

"h=\dfrac{gt^2}{2};"

"h=\dfrac{9.8\sdot (0.23\sdot10^{-6})^2}{2}=0.26\sdot10^{-12}m;"

Answer:"26. t=0.64s; L=0.64m;"

"27. t=0.23\sdot10^{-6}s; h=0.26\sdot10^{-12}m."